Write the product of (3+5i)(2+9i) in standard form
Answer by Alan3354(69443)

Write the product of (3+5i)(2+9i) in standard form
--------
Multiply like any 2 binomials.
(3+5i)(2+9i) = 6 + 37i + 45i^2 = 6 + 37i - 45
= -39 + 37i